A man firing at a distant target has 10% chance of hitting the taget in one shot.The number of times he must fire at the target to have about 50%chance of hitting the target is
a. 11
b.9
c. 7
d. 5

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Vaibhav

let he fire n number of times.

so he will miss (n-1) times and at nth time he will hit

total probability

= 1/10  + 9/10 *1/10 + 9/10 *9/10 *1/10 + .....(9/10)n-1 *1/10 = 50/100

 =1/10[ 1+ (9/10) + (9/10)2 + ............(9/10)n-1] =5/10

              or          [1- (9/10)n]/(1-9/10) =5

              0r         1-(9/10)n =1/2

                          (9/10)n  =1/2

                          nlog9/10  = log1/2

                         n =   log.5 / log.9

                            =6.58

  so he must hit at least 7 times

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin